# A Novel Hierarchical Co-Optimization Framework for Coordinated Task Scheduling and Power Dispatch in Computing Power Networks

### v2

@@ III-C2 Carbon Intensity Modeling
The environmental impact is quantified through carbon emissions. The carbon intensity of the grid is not static but depends on the real-time generation mix. The total carbon emission rate at time $t$ in scenario $\omega$, $E_{t,\omega}$, is calculated as the sum of emissions from all active generators:
@@
where $\epsilon_{g}$ is the emission factor (e.g., in tons of $CO_{2}$ per MWh) of generator $g$. For RES, $\epsilon_{g}$ is zero. This endogenous calculation is crucial, as it directly links dispatch decisions to carbon output. We also leverage real-world marginal carbon intensity data, such as that provided by WattTime, to inform the real-time DRL agent about the emissions impact of consuming an additional unit of electricity at a specific location and time.
@@ III-D Joint Optimization Problem Formulation
The overarching goal is to co-optimize the power system operation and CPN task scheduling to minimize the total expected system cost over a planning horizon $T$. The total cost comprises the operational costs of the power system and the monetized cost of carbon emissions. The problem is formulated as a large-scale, two-stage stochastic mixed-integer linear program (MILP).
@@
Key Constraints: The optimization is subject to a comprehensive set of constraints that couple the two systems:
@@
    where, $\mathcal{G}(i)$ and $\mathcal{B}(i)$ are generators and BESS at bus $i$, $P_{i,t,\omega}^{CPN}$ is the CPN power demand at that bus, and the right-hand side represents the net power flow out of the bus.
    Transmission Line Limits: The power flow $F_{ij}$ on each line $(i,j)$ must not exceed its thermal limit $F_{ij}^{max}$.
@@ IV Two-Stage Co-Optimization (TSCO) Framework
The joint optimization problem formulated in Section III is a large-scale, non-convex, mixed-integer stochastic program, which is computationally intractable to solve directly [31]. To address this challenge, we propose a Two-Stage Co-Optimization (TSCO) framework that decomposes the problem by decision timescale and complexity. The framework combines model-based optimization for long-term, system-wide planning with a model-free, AI-based approach for fast, adaptive real-time control, as shown in Fig. 2.
@@ IV-A Hierarchical Stochastic Optimization Structure
The TSCO framework decomposes the problem into two distinct stages, reflecting the natural hierarchy of power system operations:
  • Stage 1 (Day-Ahead Planning): This stage solves an SUC problem for the upcoming 24-hour horizon. It makes the ”here-and-now” decisions, which are binding across all potential future scenarios. These decisions include the commitment (on/off) status $u_{g,t}$ of conventional generators and high-level energy reservation for BESS and CPN workload classes. The objective is to minimize the total expected cost over all scenarios, setting the operational envelope for the next stage. This stage is computationally intensive but is performed only once per day.
  • Stage 2 (Real-Time Operation): This stage operates at a much faster timescale (e.g., 5-15 minute intervals) and makes ”wait-and-see” recourse decisions as uncertainty unfolds. As the actual RES generation and CPN task arrivals are revealed, this stage executes two parallel, tightly coupled processes:
@@
The interface between the two stages is designed to ensure tight coupling while maintaining computational tractability. Specifically, Stage 1 outputs three types of binding constraints to Stage 2: (1) the commitment status $u_{g,t}^{*}$ of conventional generators, which restricts the set of dispatchable units in real-time ED; (2) the energy reservation bounds $(SOC_{b,t}^{reserve,min},SOC_{b,t}^{reserve,max})$ for BESS, ensuring sufficient flexibility to accommodate real-time RES volatility; (3) the per-node computational resource reservation $C_{n,t}^{comp,reserve}$ for CPN, which limits the maximum resource utilization of each node to avoid overloading beyond the day-ahead planning envelope. In return, Stage 2 feeds back key real-time operational data to Stage 1 for iterative improvement. It includes the deviation between actual and forecasted RES generation, the CPN task completion rate and deadline adherence, and the cumulative carbon emissions up to the current time. This feedback is used to refine the scenario set $\Omega$ for the next day’s SUC problem, e.g., adjusting the probability of high RES curtailment scenarios based on real-time absorption capacity. The interface operates on a daily rolling horizon. The day-ahead plan is updated once every 24 hours, while the real-time stage fetches the latest day-ahead constraints every 5-15 minutes and reports cumulative operational data at the end of each day.
This hierarchical structure allows the framework to be both economically optimal from a long-term planning perspective and highly adaptive to short-term dynamics.
@@ IV-C DRL-based Real-Time CPN Task Scheduling
      Rolling Day-Ahead Plan Update: If the cumulative deviation between real-time CPN resource utilization and day-ahead reserve, i.e., $\sum_{t}|\sum_{k,\tau}r_{\tau}x_{k,\tau,n,t}-C_{n,t}^{comp,reserve}|$, exceeds a predefined threshold over 4 consecutive hours, the day-ahead SUC module is triggered to perform a mid-day rolling update. Then, the algorithm adjusts the remaining 24-hour reserve allocation and generator commitment status based on the latest real-time data.
@@ 10. Update BESS SOC, CPN resource status and virtual carbon queue
Additionally, the computational complexity of the TSCO framework is best analyzed by examining its two stages separately.
@@ IV-D1 Stage 1: Day-Ahead SUC
  • Subproblems: For each of the $\Omega$ scenarios, a linear program (LP) is solved. The complexity of solving an LP with modern interior-point methods is polynomial in the number of variables and constraints. Since the subproblems are independent for a given commitment schedule, they can be solved in parallel. The time taken per iteration for this step is thus equivalent to solving a single LP.
@@ V Performance Evaluation
This section presents a comprehensive empirical validation of the proposed Two-Stage Co-Optimization (TSCO) framework. A high-fidelity simulation environment is developed to assess its performance in terms of economic efficiency, environmental impact, grid stability, and CPN QoS. Additional experiments on convergence, scalability, ablation, and computational burden are conducted to further verify the framework’s engineering feasibility and core value.
@@ V-A Simulation Setup
The simulation is based on a modified IEEE 30-bus test system. It includes 6 conventional thermal generators, 4 utility-scale BESS units, and 5 large-scale renewable generation sites (3 solar, 2 wind).To ensure realism, we use real-world time-series data to model the stochastic RES generation. Solar irradiance data for locations in California is sourced from the National Renewable Energy Laboratory’s (NREL) National Solar Radiation Database (NSRDB) [^4] https://catalog.data.gov/dataset/national-solar-radiation-database-nsrdb. Wind power generation profiles for locations in Germany are obtained from the ENTSO-E Transparency Platform [^5] https://www.entsoe.eu/data/power-stats/. These datasets are used to generate 100 distinct 24-hour scenarios for the SUC problem.
@@ V-B Comparison Schemes
  • Cost-Only Optimizer (CO-Opt) [35]: This baseline uses the same two-stage optimization architecture as TSCO but with the carbon price set to zero. It represents the current industry-standard approach of economic dispatch, which focuses exclusively on minimizing direct economic costs without consideration for environmental impact.
  • Renewable-Greedy Scheduler (RG-Sched) [36]: This is a heuristic-based CPN scheduling approach where tasks are always dispatched to the CPN node with the highest instantaneous RES power availability. The power system is not co-optimized; it simply reacts to the resulting CPN load profile. This baseline tests the efficacy of a simple “follow the renewables” strategy that ignores grid constraints and economic signals.
  • Decoupled Framework (DC-Frame) [37]: It is the state-of-the-art in carbon-aware computing, where the power system and CPN are optimized separately. The power system operation is optimized first to generate a fixed 24-hour profile of electricity prices and carbon intensities. Subsequently, the CPN scheduler optimizes its task scheduling based on these static, pre-computed signals. This approach is carbon-aware but lacks the tight, real-time feedback loop of our co-optimization framework.
@@ V-C1 Baseline Performance Comparison
In this scenario, all four methods were simulated over one week (168 hours) with the carbon price set at a representative value of $\mathdollar 50$/ton. The mean and standard deviation are present in Fig. 3.
TSCO simultaneously achieves the lowest total operational cost and the lowest carbon emissions. The CO-Opt baseline, being carbon-agnostic, minimizes only direct fuel costs, resulting in $41.5\%$ higher emissions and $17.5\%$ higher total costs once the carbon price is factored in. This is because it relies heavily on the cheapest available thermal generators, regardless of their carbon intensity. The RG-Sched heuristic, while intuitive, performs poorly across the board. By myopically chasing renewables, it ignores grid congestion and the economic cost of dispatching thermal generators to support its decisions, leading to the highest operational cost and only modest emission reductions. The DC-Frame performs better than the naive baselines but is still significantly outperformed by TSCO. Its reliance on static, day-ahead signals prevents it from adapting to real-time deviations between forecasted and actual grid conditions, leading to $12.7\%$ higher costs and $16.2\%$ higher emissions. A key finding is TSCO’s ability to significantly reduce RES curtailment. By treating the CPN as a flexible load that can absorb surplus renewable generation in real-time, TSCO reduces curtailment by over $60\%$ compared to DC-Frame and CO-Opt. This demonstrates the value of co-optimization in turning a major energy consumer into a valuable grid-stabilizing asset.
@@ V-C2 Sensitivity to Carbon Price
To analyze the trade-off between economic and environmental objectives, we varied the carbon price $\lambda_{CO_{2}}$ from $\mathdollar 25$/ton to $\mathdollar 150$/ton. This analysis focuses on the TSCO and DC-Frame methods, as the other two baselines are insensitive to carbon price by design. The comprehensive results are presented in Fig. 4.
The TSCO framework consistently achieves a better trade-off. For any given carbon emission level, it operates at a lower cost than the DC-Frame. As the carbon price increases, both methods are incentivized to reduce emissions, but TSCO does so more efficiently. This superior performance is a direct result of the real-time feedback loop between the grid and the CPN, which allows the system to find more efficient operating points. The RG-Sched method is shown as a fixed point, as their operational strategy does not adapt to the carbon price.
@@
It also shows how RES curtailment is affected by the carbon price. For both carbon-aware methods, increasing the price of carbon incentivizes greater utilization of zero-emission renewable energy, thus reducing curtailment. However, the TSCO framework’s ability to react to real-time conditions allows it to absorb significantly more renewable energy across all price points, maintaining a curtailment level that is less than half that of the DC-Frame. This again underscores the value of tight system integration for maximizing the use of clean energy resources.
@@ V-D Convergence Comparison
To verify the efficiency and stability of the Benders decomposition in the day-ahead SUC stage, we compare two methods under different load intensities, including Direct-MILP without decomposition and Improved-Benders in [32]. The results are summarized in Table II. Among them, the mid-range workload represents an average daily task arrival rate of $800$ jobs per day with a total computational workload of $5\times 10^{15}$ FLOPs, and an average resource requirement of $4$ CPU cores and 2 GPU cores or equivalent TPU resources per job. And the low and high loads are $0.8$ times and $1.2$ times the medium load, respectively.
@@ V-G Analysis of Computational Burden
To verify the engineering feasibility of the TSCO framework, we quantify the computational time of each stage under a standard server environment (CPU: Intel Xeon 8375C 3.0GHz; GPU: NVIDIA A100 40GB; Memory: 256GB). The results are shown in Table IV.
@@ VI Conclusion
This paper has addressed the critical and intertwined challenges of rising energy consumption in CPNs and the increasing instability of power grids due to high renewable energy penetration. We have argued that treating these two complex systems in a decoupled manner leads to suboptimal outcomes, characterized by high costs, significant carbon emissions, and wasted renewable energy. To overcome these limitations, a novel TSCO framework was proposed, designed to synergistically manage power system operations and CPN task scheduling. The results clearly demonstrate that by enabling the CPN to act as an active, flexible participant in grid operations, significant benefits can be realized. Compared to a decoupled approach, our integrated TSCO framework reduced total operational costs and carbon emissions. Most notably, it slashed renewable energy curtailment by over $60\%$ compared to conventional cost-only optimization, all while maintaining a high job success rate of over $98.5\%$ for computational tasks.

### v3

@@ III-C2 Carbon Intensity Modeling
The environmental impact is quantified through carbon emissions. The carbon intensity is not static but depends on the real-time generation mix. The total carbon emission rate at time $t$ in scenario $\omega$, $E_{t,\omega}$, is calculated as the sum of emissions from all active generators:
@@
where $\epsilon_{g}$ is the emission factor (e.g., in tons of $CO_{2}$ per MWh) of generator $g$. For RES, $\epsilon_{g}$ is zero. This endogenous calculation is crucial, as it directly links dispatch decisions to carbon output. We also leverage real-world marginal carbon intensity data, such as that provided by WattTime [^2] https://watttime.org/data-science/data-signals/, which informs the real-time DRL agent about the emission impact resulting from an additional unit of electricity consumption at a specific location and time.
@@ III-D Joint Optimization Problem Formulation
The overarching goal is to co-optimize the power system operation and CPN task scheduling to minimize the total expected system cost over a planning horizon $T$. It comprises the operational costs of the power system and the monetized cost of carbon emissions. It is formulated as a large-scale, two-stage stochastic mixed-integer linear program (MILP).
@@
The optimization is subject to a comprehensive set of constraints that couple the two systems:
@@
    where $\mathcal{G}(i)$ and $\mathcal{B}(i)$ are generators and BESS at bus $i$. $P_{i,t,\omega}^{CPN}$ is the CPN power demand. The right side represents the net power flow out of the bus.
    Transmission Line Constraints: The power flow $F_{ij}$ on each line $(i,j)$ must not exceed its thermal limit $F_{ij}^{max}$.
@@ IV Two-Stage Co-Optimization (TSCO) Framework
The joint optimization problem formulated in Section III is a large-scale, non-convex, mixed-integer stochastic program, whose direct solution is computationally prohibitive at realistic scales [32]. Our TSCO framework combines model-based optimization for long-term, system-wide planning with a model-free, AI-based approach for fast, adaptive real-time control, as shown in Fig. 2.
@@ IV-A Hierarchical Stochastic Optimization Structure
The TSCO framework decomposes the problem into two distinct parts, reflecting the power system’s natural hierarchy.
  • Stage 1 (Day-Ahead Planning): This stage solves an SUC problem for the upcoming 24-hour horizon. It makes the “here-and-now” decisions, which are binding across all potential future scenarios. These decisions include the commitment (on/off) status $u_{g,t}$ of conventional generators and high-level energy reservation for BESS and CPN workload classes. The objective is to minimize the total expected cost over all scenarios, setting the operational envelope for the next stage. This stage is computationally intensive but is performed only once per day.
  • Stage 2 (Real-Time Operation): This stage operates at a much faster timescale (e.g., 5-15 minute intervals) and makes “wait-and-see” recourse decisions as uncertainty unfolds. As the actual RES generation and CPN task arrivals are revealed, this stage executes two parallel, tightly coupled processes:
@@
The interface between the two stages is designed to ensure tight coupling while maintaining computational tractability. Stage 1 outputs three types of binding constraints to Stage 2: (1) the commitment status $u_{g,t}^{*}$ of conventional generators, which restricts the set of dispatchable units in real-time ED; (2) the energy reservation bounds $(SOC_{b,t}^{reserve,min},SOC_{b,t}^{reserve,max})$ for BESS, ensuring sufficient flexibility to accommodate real-time RES volatility; (3) the per-node computational resource reservation $C_{n,t}^{comp,reserve}$ for CPN, which limits the maximum resource utilization of each node to avoid overloading beyond the day-ahead planning envelope. In return, Stage 2 feeds back key real-time operational data to Stage 1 for iterative improvement. It includes the deviation between actual and forecasted RES generation, the CPN task completion rate and deadline adherence, and the cumulative carbon emissions up to the current time. This feedback is used to refine the scenario set $\Omega$ for the next day’s SUC problem.The day-ahead plan is updated once every 24 hours, while the real-time stage fetches the latest day-ahead constraints every 5-15 minutes and reports cumulative operational data at the end of each day.
@@ IV-C DRL-based Real-Time CPN Task Scheduling
      Rolling Day-Ahead Plan Update: If the cumulative deviation between real-time CPN resource utilization and day-ahead reserve, i.e., $\sum_{t}|\sum_{k,\tau}r_{\tau}x_{k,\tau,n,t}-C_{n,t}^{comp,reserve}|$, exceeds a predefined threshold over 4 consecutive hours, the day-ahead SUC module is triggered to perform a mid-day rolling update. Then, the algorithm adjusts the remaining 24-hour reserve allocation and generator commitment status based on the real-time data.
@@ 10. Update BESS SOC, CPN resource status and virtual carbon queue
Additionally, the computational complexity of the TSCO framework is best analyzed by examining its two stages.
@@ IV-D1 Stage 1: Day-Ahead SUC
  • Subproblems: For each $\Omega$ scenario, a linear program (LP) is solved. The complexity of solving an LP with modern interior-point methods is polynomial in the number of variables and constraints. Since the subproblems are independent for a given commitment schedule, they can be solved in parallel. The time taken per iteration for this step is thus equivalent to solving a single LP.
@@ V Performance Evaluation
A high-fidelity simulation environment is developed to assess its performance in terms of economic efficiency, environmental impact, grid stability, and CPN QoS. Additional experiments on convergence, scalability, ablation, and computational burden are conducted to further verify the framework’s engineering feasibility and core value.
@@ V-A Simulation Setup
The simulation is based on a modified IEEE 30-bus test system [^3] https://icseg.iti.illinois.edu/ieee-30-bus-system/. It includes 6 conventional thermal generators, 4 utility-scale BESS units, and 5 large-scale renewable generation sites (3 solar, 2 wind).To ensure realism, we use real-world time-series data to model the stochastic RES generation. Solar irradiance data for locations in California is sourced from the National Renewable Energy Laboratory’s (NREL) National Solar Radiation Database (NSRDB) [^4] https://catalog.data.gov/dataset/national-solar-radiation-database-nsrdb. Wind power generation profiles for locations in Germany are obtained from the ENTSO-E Transparency Platform [^5] https://www.entsoe.eu/data/power-stats/. These datasets are used to generate 100 distinct 24-hour scenarios for the SUC problem.
@@ V-B Comparison Schemes
  • Cost-Only Optimizer (CO-Opt) [36]: This baseline uses the same two-stage optimization architecture as TSCO but with the carbon price set to zero. It represents the current industry-standard approach of economic dispatch.
  • Renewable-Greedy Scheduler (RG-Sched) [37]: This is a heuristic-based CPN scheduling approach where tasks are always dispatched to the CPN node with the highest instantaneous RES power availability.
  • Decoupled Framework (DC-Frame) [17]: It is the state-of-the-art in carbon-aware computing, where the power system and CPN are optimized separately. The power system operation is optimized first to generate a fixed 24-hour profile of electricity prices and carbon intensities. Subsequently, the CPN scheduler optimizes its task scheduling based on these static, pre-computed signals.
  • DRL-based Spatiotemporal Scheduler (ST-Sched) [15], [16]: This baseline adopts a customized DRL approach for intra-CPN fine-grained scheduling. It features a triple-selection action space to optimize energy consumption, carbon emissions, and task delay while ensuring load balancing.
@@ V-C1 Baseline Performance Comparison
In this scenario, all five methods were simulated over one week (168 hours) with the carbon price set at a representative value of $\mathdollar 50$/ton. The mean and standard deviation are present in Fig. 3.
TSCO simultaneously achieves the lowest total operational cost and the lowest carbon emissions. The CO-Opt baseline, being carbon-agnostic, minimizes only direct fuel costs, resulting in $41.5\%$ higher emissions and $17.5\%$ higher total costs once the carbon price is factored in. This is because it relies heavily on the cheapest available thermal generators, regardless of their carbon intensity. The RG-Sched heuristic, while intuitive, performs poorly across the board. By myopically chasing renewables, it ignores grid congestion and the economic cost of dispatching thermal generators to support its decisions, leading to the highest operational cost and only modest emission reductions. The DC-Frame performs better than the naive baselines but is still significantly outperformed by TSCO. Its reliance on static, day-ahead signals prevents it from adapting to real-time deviations between forecasted and actual grid conditions, leading to $12.7\%$ higher costs and $16.2\%$ higher emissions. The ST-Sched focuses on intra-CPN fine-grained scheduling via customized DRL. It demonstrates competitive performance relative to DC-Frame and CO-Opt but falls short of TSCO. It achieves $8.4\%$ higher operational costs and $4.4\%$ higher carbon emissions than TSCO. Its lack of cross-domain coordination with the power system prevents it from leveraging real-time grid dynamics (e.g., surplus renewable energy) to optimize overall efficiency. A key finding is TSCO’s ability to significantly reduce RES curtailment. By treating the CPN as a flexible load that can absorb surplus renewable generation in real-time, TSCO reduces curtailment by over $60\%$ compared to DC-Frame and CO-Opt, and by $37.1\%$ compared to ST-Sched. This demonstrates the value of co-optimization in turning a major energy consumer into a valuable grid-stabilizing asset.
@@ V-C2 Sensitivity to Carbon Price
To analyze the trade-off between economic and environmental objectives, we varied the carbon price $\lambda_{CO_{2}}$ from $25$ to 150150/ton. This analysis focuses on the TSCO, DC-Frame, and ST-Sched methods, as the other two baselines are insensitive to carbon price by design. The comprehensive results are presented in Fig. 4.
TSCO achieves a superior economic-environmental trade-off across all metrics at every carbon price point. All three carbon-aware methods cut emissions with rising carbon prices, but TSCO does so with the highest efficiency at $19.3\%$. Economically, TSCO also demonstrates optimal performance. Its operational cost rises by the smallest margin with increasing carbon prices, remaining approximately $7\%$ lower than ST-Sched and $14\%$ lower than DC-Frame across all price points. ST-Sched’s single-domain intra-CPN DRL optimization delivers moderate cost and emission improvements over DC-Frame but falls short of TSCO. It lacks cross-domain coordination to leverage grid-side low-carbon, low-cost resources. DC-Frame’s decoupled design, reliant on static day-ahead grid signals, fails to adapt to real-time deviations. It leads to the highest operational costs and the least efficient emission reduction among them. For RES curtailment, rising carbon prices drive greater renewable utilization for all methods, with TSCO recording the largest reduction at $14.3\%$. Its curtailment levels are $37.1\%$ lower than ST-Sched and less than half that of DC-Frame. This underscores TSCO’s unique ability to treat the CPN as a flexible load to absorb surplus grid-side renewable energy.
@@
Notably, TSCO’s environmental and economic gains do not compromise CPN service quality. It retains the highest job success rate and lowest average tardiness across all carbon prices. ST-Sched also delivers strong QoS performance due to its refined intra-CPN load balancing. While DC-Frame experiences noticeable QoS degradation with rising carbon prices due to its rigid precomputed grid signals failing to account for real-time CPN congestion.
@@ V-D Convergence Comparison
To verify the efficiency and stability of the Benders decomposition in the day-ahead SUC stage, we compare two methods under different load intensities, including Direct-MILP without decomposition and Improved-Benders [33]. The results are summarized in Table II. Among them, the mid-range workload represents an average daily task arrival rate of $800$ jobs per day with a total computational workload of $5\times 10^{15}$ FLOPs, and an average resource requirement of $4$ CPU cores and $2$ GPU cores or equivalent TPU resources per job. The low and high loads are $0.8$ times and $1.2$ times the medium load, respectively.
@@ V-G Analysis of Computational Burden
To verify the feasibility of the TSCO framework, we quantify the computational time of each stage under a standard server environment (CPU: Intel Xeon 8375C 3.0GHz; GPU: NVIDIA A100 40GB; Memory: 256GB). The results are shown in Table IV.
@@ VI Conclusion
This paper addresses the intertwined challenges of growing energy consumption in CPNs and power grid instability caused by high RES penetration. These two issues are critical for service computing sustainability and QoS. To overcome these limitations, we propose a novel TSCO framework that synergistically coordinates CPN task scheduling and power system dispatch. By decomposing the complex problem into a day-ahead stochastic unit commitment stage and a real-time operational stage, the framework aligns service QoS requirements with grid sustainability goals. Extensive simulations validate the TSCO framework’s superiority over baselines. It achieves a $16.2\%$ cut in carbon emissions and a $12.7\%$ reduction in operational costs, while cutting RES curtailment by more than $60\%$. Notably, these environmental and economic gains are achieved without compromising service reliability. The framework sustains a $98.5\%$ task success rate and limits average task tardiness to $12.3$s. This work advances cross-domain service optimization in CPNs, laying the groundwork for efficient, low-carbon service computing.